### v1

# TimeSieve: Extracting Temporal Dynamics through Information Bottleneck

Abstract.
Time series forecasting has become an increasingly popular research area due to its critical applications in various real-world domains such as traffic management, weather prediction, and financial analysis. Despite significant advancements, existing models face notable challenges, including the necessity of manual hyperparameter tuning for different datasets, and difficulty in effectively distinguishing signal from redundant features in data characterized by strong seasonality. These issues hinder the generalization and practical application of time series forecasting models. To solve this issues, we propose an innovative time series forecasting model - TimeSieve designed to address these challenges. Our approach employs wavelet transforms to preprocess time series data, effectively capturing multi-scale features without the need for additional parameters or manual hyperparameter tuning. Additionally, we introduce the information bottleneck theory that filters out redundant features from both detail and approximation coefficients, retaining only the most predictive information. This combination reduces significantly improves the model’s accuracy. Extensive experiments demonstrate that our model outperforms existing state-of-the-art methods on 70% of the datasets, achieving higher predictive accuracy and better generalization across diverse datasets. Our results validate the effectiveness of our approach in addressing the key challenges in time series forecasting, paving the way for more reliable and efficient predictive models in practical applications. The code for our model is available at https://github.com/xll0328/TimeSieve.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
[…]
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
1. Introduction
The field of time series forecasting has experienced significant advancements and found widespread applications in various real-world engineering tasks, such as traffic management (Lana et al., 2018), pandemic spread (Kumar and Susan, 2020), and financial analysis (Sezer et al., 2020). From early models like GRU (Dey and Salem, 2017) and LSTM (Greff et al., 2016) to the more recent Transformer-based architectures (Liu et al., 2023b; Wu et al., 2021), a plethora of sophisticated models have emerged. For instance, TimesNet (Wu et al., 2023) and FreTS (Yi et al., 2023) approach data processing from a frequency domain perspective, while multi-scale models like Pyraformer (Liu et al., 2022) and Scaleformer (Shabani et al., 2023) leverage intricate multi-scale and cross-dimensional dependency structures to significantly enhance performance on complex forecasting tasks.
Howerver, existing models often require the introduction of additional parameters to capture multi-scale features, which increases computational complexity. Moreover, these models necessitate manual tuning of hyperparameters for different datasets, further complicating their application and limiting their generalization across diverse scenarios. More importantly, many models, including Autoformer (Wu et al., 2021) , struggle to effectively distinguish between signal and redundant features when dealing with data characterized by strong seasonality. Redundant features not only interferes with the learning process, leading to increased errors, but also obscures the true patterns within the data, making the predictions unreliable. This issue is particularly pronounced when traditional autocorrelation mechanisms break down due to redundant features interference, significantly diminishing the practical value of these models.
To address these challenges, we proposes an innovative time series forecasting model designed to significantly improve predictive accuracy and broaden applicability. First, we employ wavelet transforms to preprocess the time series data using the Wavelet Decomposition Block (WDB) and the Wavelet Reconstruction Block (WRB). Wavelet transforms efficiently capture information from multiple scales, but this process may also introduce redundant features. To address this, we integrate the Information Filtering and Compression Block (IFCB), an information bottleneck module that compresses and extracts the most relevant features by filtering out redundant features from both detail and approximation coefficients. These components work together, significantly enhancing the model’s accuracy.
Our main contributions can be summarized as follows:
  • We introduce WDB and WRB that efficiently capture multi-scale features in time series data without requiring manual hyperparameter tuning or the introduction of additional parameters.
  • We utilize an information-theoretic approach by incorporating an IFCB to filter out redundant features from the data. This ensures that only the most predictive features are retained, enhancing the model’s accuracy.
  • By integrating wavelet transforms for comprehensive feature extraction and the information bottleneck block for redundant features, our model achieves state-of-the-art performance on 70% of the datasets tested, demonstrating superior predictive accuracy and generalization across diverse scenarios.
The rest of this paper is organized as follows. Section 2 reviews related work in the field of time series forecasting. In Section 3, we provide a detailed description of our model implementation. The experimental results and analysis are discussed in Section 4. Finally, Section 5 concludes the paper with a summary of our findings and suggestions for future research.
2. Related Work
2.1. Time Series Forecasting
In recent years, the field of time series forecasting has advanced rapidly, giving rise to many efficient models. These models predominantly rely on multilayer perceptron (MLP) approaches like MSD-Mixer (Zhong et al., 2024), TSMixer (Chen et al., 2023),DLinear (Zeng et al., 2022)and FreTs (Yi et al., 2023), along with Transformer-based techniques such as PatchTST (Nie et al., 2023), Autoformer, and Informer (Zhou et al., 2021), to perform predictions using sophisticated data transformation and learning mechanisms.
Time series forecasting can be defined as follows: Given a historical sequence of observations $\ X=\{x_{1},x_{2},\ldots,x_{T}\}$, where $T$ is the lookback window length, the goal is to predict future values $\ Y=\{y_{T+1},y_{T+2},\ldots,y_{T+H}\}$, where $H$ is the forecast window length.
(1) $Y=F(X)$
where $F$ is the predictive model mapping the input sequence $X$ to the output sequence $Y$. However, time series data often exhibit complex, multi-scale characteristics. Conventional feature extraction methods, like standard pooling, require additional parameter adjustments or rely on manually designed pooling sizes, increasing model complexity and limiting flexibility.
Wavelet transform, a classical signal processing technique (Farge, 1992), excels in multi-scale analysis. It decomposes signals into different frequency components, aiding in the understanding and prediction of intricate patterns.
We proposes a novel approach that combines wavelet transform with modern machine learning techniques for time series forecasting. The wavelet transform preprocesses the input data, decomposing it into detail coefficients and approximation coefficients at different scales. We believe this method effectively captures comprehensive features from the time series data.
2.2. Information Bottleneck Theory
In recent years, with the growing intersection of information theory and deep learning research, many researchers have begun analyzing and optimizing neural networks from the perspective of information theory. In this trend, Information Bottleneck (IB) theory has gained particular attention (Tishby and Zaslavsky, 2015; Shwartz-Ziv and Tishby, 2017), finding wide application in emerging fields such as graph neural networks (Dai et al., 2023), multimodal learning (Mai et al., 2022), and recently in time series imputation (Choi and Lee, 2024). IB theory helps improve model performance by optimizing information flow and processing (Tishby and Zaslavsky, 2015).
We believe that wavelet transformation can effectively decompose time series data into components that contain valuable information. However, these components also include a significant amount of redundant information. If not properly addressed, this redundant information can interfere with the model’s learning performance and predictive accuracy. From the perspective of information theory, IB theory provides a mechanism to effectively retain critical information related to target variables while filtering out irrelevant or redundant information in the model’s intermediate representations. This approach enhances the model’s ability to capture key features. Furthermore, it strikes a delicate balance between compressing input information and maintaining predictive accuracy, ensuring that the model’s output remains precise and concise by:
(2) $\max I(y;z)$
$\min I(x;z)$
where $I(y;z)$ represents the mutual information between the output $y$ and the intermediate representation $z$, and $I(x;z)$ represents the mutual information between the input $x$and the intermediate representation $z$. Maximizing $I(y;z)$ ensures that the most relevant information for predicting the output is retained, while minimizing $I(x;z)$ helps to compress the input information, filtering out redundant features.
Therefore, this work seeks to apply information-theoretic methods to time series forecasting. Unlike prior applications of the IB theory in deep learning, we do not simply apply it directly. Instead, we integrate it into the overall framework to specifically handle the componets obtained from the wavelet transformation. In this way, IB theory can effectively filter out redundant features from these coefficients while retaining key information, ultimately enhancing the model’s predictive performance.
3. TimeSieve
Figure 1. Overview of the TimeSieve Framework. The upper part illustrates the overall architecture of TimeSieve, which consists of the Wavelet Decomposition Block (WDB), Wavelet Reconstruction Block (WRB), Information Fusion and Compression Block (IFCB), and the prediction. The input time series data, $T\times C$, is decomposed by WDB into coefficients $cA$ and $cD$, which are processed by IFCB. The WRB then reconstructs the data, followed by the predictor generating the corresponding forecast steps. The lower part provides a detailed visualization of the data flow within WDB and IFCB, with further visualization of IFCB available in Figure 2.
In our study, we propose a novel time series forecasting model named TimeSieve. This model integrates the Information Filtering and Compression Block (IFCB) with wavelet transform technology to enhance the accuracy of time series predictions.
To effectively capture comprehensive feature information, we employ the Wavelet Decomposition Block (WDB). The WDB decomposes time series data into different frequency components, effectively extracting multi-scale information. Specifically, the wavelet transform can be represented by the following equations:
(3) $[cA,cD]=WDB(X)$
where the approximation coefficients $cA$ representing the low-frequency trends, and the detail coefficients $cD$ representing the high-frequency details.
However, the extracted multi-scale information may contain redundant features, which can adversely affect the model’s learning performance and predictive accuracy. To address this issue, we introduce the Information Filtering and Compression Block (IFCB) to filter the information. Additionally, we employ residual connections to ensure that valuable information is not lost during the filtering process. The IFCB optimizes the information flow, retaining critical information while filtering out irrelevant or redundant information. The equations for the Information Bottleneck module are as follows:
(4) $c\hat{A}=IFCB(cA)+cA$
(5) $c\hat{D}=IFCB(cD)+cD$
where $c\hat{A}$ and $c\hat{D}$ represents the filtered coefficients $cA$ and $cD$. After filtering with the IFCB, we apply the Wavelet Reconstruction Block (WRB) to reconstruct the processed data back into the time domain. This step ensures that the features at different scales are fully utilized. Finally, we use a Multi-Layer Perceptron (MLP) as the prediction to make the final forecast on the processed time series data. The equation for the prediction step is as follows:
(6) $Y=MLP(WRB([c\hat{A},c\hat{D}]))$
The overall architecture of the TimeSieve model is illustrated in Figure 1. The TimeSieve model leverages the combination of wavelet transform and IFCB to effectively handle the multi-scale characteristics present in time series data, thereby improving the model’s predictive performance. This approach is motivated by the need to optimize information flow and enhance feature extraction, ensuring that our model can make accurate and reliable predictions across various applications.
The above is an overall description of TimeSieve. We will provide a detailed introduction to WDB, WRB, and IFCB in the following sections.
3.1. WDB and WRB
To effectively capture comprehensive feature information, we employ wavelet transformation to decompose time series data inspired by (Meyer et al., 2023; Nason and Wei, 2024), allowing us to analyze the data across multiple scales. Wavelet transformation allows us to break down the data into various frequency components, enabling us to analyze and extract features across multiple temporal scales. This multi-scale analysis is crucial for identifying patterns and trends that occur at different time resolutions.
WDB. For the input time series $x(t)\in\mathbb{R}^{T\times C}$, where $T$ represents the length of the lookback window and $C$ denotes the number of variates, the wavelet transform decomposes the series into approximation coefficients $cA$ and detail coefficients $cD$. The approximation coefficients $cA$ capture the general trend or smooth component of the series, representing low-frequency information. In contrast, the detail coefficients $cD$ capture the fine-grained, high-frequency information, highlighting local fluctuations and details within the series. This decomposition can be expressed as:
[…]
where $\phi(t)$ and $\psi(t)$ are the scaling function and wavelet function, respectively, defined as follows:
(9) $\phi(x)=\sum_{k=0}^{N-1}a_{k}\phi(2u-k),$
(10) $\psi(x)=\sum_{k=0}^{M-1}b_{k}\phi(2u-k),$
In the case of db1 wavelet transform, $\phi(t)$ represents the low-pass filter, capturing the smooth, approximate part of the signal, and $\psi(t)$ represents the high-pass filter, capturing the detailed, high-frequency part of the signal. Here, $a_{k}$ and $b_{k}$ are the filter coefficients for the scaling and wavelet functions, respectively, $u$ is the variable over which the functions are defined, and $k$ is the index of the coefficients.
This decomposition allows the model to capture different characteristics of the data at various levels, thereby optimizing the understanding and prediction of time series dynamics, particularly in identifying local features within the series.
WRB. After the wavelet transformation, the detail and approximation coefficients are processed through IFCB to preserve the most predictive information. Subsequently, the inverse wavelet transform is applied to reconstruct the time series from the processed coefficients:
(11) $\hat{x}(t)=\sum c\hat{A}\phi(t)+\sum c\hat{D}\psi(t)$
where $c\hat{A}$ and $c\hat{D}$ are the filtered approximation and detail coefficients, respectively. Finally, the reconstructed time series is passed through a MLP for final prediction.
3.2. IFCB
Motivated by the need to effectively filter out redundant features while retaining critical information, we introduce the Information Filtering and Compression Block (IFCB), inspired by the information bottleneck principle (Tishby and Zaslavsky, 2015; Shwartz-Ziv and Tishby, 2017; Liu et al., 2024). In this section, we process the detail coefficients $cD$ and approximation coefficients $cA$ from the wavelet transform of the time series data, for convenience in subsequent explanations, we define $cI\in(cA,cD)$ to collectively represent $cA$ and $cD$. Additionally, We use $c\hat{I}$, where $I$ is in the range $[A,D]$. to represent $c\hat{A}$ and $c\hat{D}$, the filtered versions of $cA$ and $cD$. Initially, we define the input to the IFCB as $cI\in\mathbb{R}^{T/2\times C}$ and the output to the IFCB as $c\hat{I}\in\mathbb{R}^{T/2\times C}$.
In our model, both $cI$ and $c\hat{I}$ are treated as random variables with a joint distribution $p(i,\hat{i})$. The mutual information between them, denoted as $I(cI;c\hat{I})$, quantifies the amount of information shared by $cI$ and $c\hat{I}$ (More details you can see in Appendix B). It is defined as:
(12) $1(cI;c\hat{I})$ $=D_{kL}[p(i,\hat{i})\|p(i)p(\hat{i})]$
We define the intermediate hidden layer random variable $Z$. Following the principles of a Markov chain where $cI\rightarrow Z\rightarrow c\hat{I}$, the mutual information satisfies $I(cI;Z)\geq I(cI;c\hat{I})$ (More details see in Appendix A). Although there is an inevitable information loss in the transmission to the middle layer $Z$, the key is that $Z$ retains the most critical information for predicting $c\hat{I}$ while filtering out irrelevant features.
This allows us to define the objective of IFCB, which is to maximize the mutual information between $Z$ and $c\hat{I}$ while minimizing the information between $Z$ and $cI$.
(13) $\min\{I(cI;Z)-\beta I(Z;c\hat{I})\}$
where $\beta$ is a trade-off parameter that determines the weight of the minimal information constraint in the optimization process.
The optimization is carried out using a deep neural network, which is designed to minimize the above objective by adjusting the network weights and biases through backpropagation and suitable regularization techniques.
Figure 2. Overview of IFCB: This module consists of a reparameterization step and deep neural networks. Initially, the standard deviation and mean ($\mu_{z},\Sigma_{z}$) of the data are computed. The reparameterization is then performed using the equation $z=\mu_{z}+\Sigma_{z}\times\epsilon$. The input coefficients $cI$ are processed through the first neural network to obtain $\mu_{z}$ and $\Sigma_{z}$. The reparameterized $z$ is then fed into the second neural network to produce the reconstructed coefficients $\hat{cI}$.
Optimization is then performed using a deep neural network. Assuming $p(z|i)$ follows a Gaussian distribution, we can derive the following:
(14) $p(z|i)$ $=\mathcal{N}(\mu(i;\theta_{\mu}),\Sigma(i;\theta_{\Sigma}))$
$=\mathcal{N}(\mu_{z},\Sigma_{z})$
where $\theta_{\mu}$ and $\theta_{\Sigma}$ are the parameters of the networks that are optimized during training.
Parameter updates pose a challenge due to the stochastic nature of the gradient calculations. To address this, the reparameterization trick is introduced. The formula is given as:
(15) $z=\mu_{z}+\Sigma_{z}\times\epsilon$
where $\epsilon$ is sampled from a standard Gaussian distribution.
Given that the study focuses on a regression task, we define the conditional probability $q(\hat{i}|z)$ as follows:
(16) $q(\hat{i}|z)=e^{-||\hat{i}-D(z;\theta_{c})||}+C$
where $D$ is a decoder function parameterized by $\theta_{c}$, and $C$ is a constant. This formulation measures the fit of the model’s predictions by capturing the deviation of the predicted output $\hat{i}$ from the output of the decoder.
We now specify the model’s loss function. This function is crucial for training the model to minimize prediction errors and optimize the distribution parameters effectively. The loss is composed of the original loss component and the IFCB loss, as shown below:
(17) $\mathcal{L}$ $=\mathcal{L}_{o}+\mathcal{L}_{IB}$
$=\mathcal{L}_{o}+D_{KL}[\mathcal{N}(\mu_{z},\Sigma_{z})\,||\,\mathcal{N}(0,I)]+D_{KL}[p(z)\,||\,p(z|i)]$
where $\mathcal{L}_{o}$ is the original prediction task loss, typically representing the error in regression predictions, and $\mathcal{L}_{IB}$is the IFCB loss.
The overall architecture of the IFCB is illustrated in Figure 2.
Algorithm 1 Pseudo code of TimeSieve
0:  Time series data $X=\{x_{1},x_{2},\ldots,x_{T}\}$, true labels $Y=\{y_{T+1},y_{T+2},\ldots,y_{T+H}\}$
0:  Predicted output $\hat{Y}$
1:  Initialization: Initialize the model parameters, learning rate $\eta$, and total training epochs $E$
2:  Training:
3:  for $epoch\leftarrow 1$ to $E$ do
4:     for $i\leftarrow 1$ to $N$ do
[…]
6:        for $I\in\{A,D\}$ do
7:           $c\hat{I},\mu_{z},\Sigma_{z}\leftarrow\text{IFCB}(cI)$
8:        end for
9:        $X_{\text{reconstructed}}\leftarrow$ WRB($c\hat{A},c\hat{D}$)
10:        $\hat{Y}_{i}\leftarrow\text{MLP}(X_{\text{reconstructed}})$
11:        $\mathcal{L}_{o}\leftarrow\text{LossFunction}(\hat{Y}_{i},Y_{i})$
12:        for $I\in\{A,D\}$ do
13:           $\mathcal{L}_{IB_{I}}\leftarrow D_{KL}[\mathcal{N}(\mu_{zI},\Sigma_{zI})\,||\,\mathcal{N}(0,I)]+D_{KL}[p(z_{I})\,||\,p(z_{I}|i)]$
14:        end for
15:        $\mathcal{L}_{IB}\leftarrow\mathcal{L}_{IB_{A}}+\mathcal{L}_{IB_{D}}$
16:        $\mathcal{L}\leftarrow\mathcal{L}_{o}+\mathcal{L}_{IB}$
17:        $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
18:     end for
19:  end for
20:  Inference:
21:  for $i\leftarrow 1$ to $N$ do
[…]
23:     for $I\in\{A,D\}$ do
24:        $c\hat{I},\mu_{z},\Sigma_{z}\leftarrow\text{IFCB}(cI)$
25:     end for
26:     $X_{\text{reconstructed}}\leftarrow$ WRB($c\hat{A},c\hat{D}$)
27:     $\hat{Y}_{i}\leftarrow\text{MLP}(X_{\text{reconstructed}})$
28:  end for
29:  return  final predictions $\hat{Y}$
4. Experiments
4.1. Datasets and Baselines
We conduct comprehensive experiments to assess the performance of TimeSieve. For multivariate forecasting, we utilize a variety of real-world benchmarks, including ETT (Zhou et al., 2021), Exchange (Lai et al., 2018), Electricity(ECL) (Lin et al., 2023), and National Illness(ILI) (Zeng et al., 2022). Notably, instead of using a fixed lookback window length, we follow the approach of Koopa (Liu et al., 2023a) by setting the lookback window length $T$ to twice the forecast window length (i.e., $T=2H$). This approach is based on the fact that in real-world scenarios, historical observations are readily accessible, and utilizing a larger amount of observed data can enhance the performance of deep models, particularly as the forecast horizon extends (Liu et al., 2023a).
We extensively compare TimeSieve with state-of-the-art deep forecasting models, including Transformer-based models such as Nonstationary-Transformer (NSTformer) (Liu et al., 2023b) and Autoformer (Wu et al., 2021); MLP-based models such as LightTS (Campos et al., 2023), and Koopman operator-based models such as Koopa (Liu et al., 2023a). See the Appendix C for a more detailed introduction.
Table 1. Forecasting results with different forecast lengths H $\in$ $\{24,36,48,60\}$ for ILI and H $\in$ $\{48,96,144,192\}$ for others. We set the lookback length T = 2H. Red indicates the best result, while Blue refers to the second best.
Models TimeSieve Koopa NSTformer LingtTS Autoformer
[…]
ETTh1 48 0.361 0.341 0.385 0.364 0.465 0.614 0.406 0.404 0.432 0.678
96 0.384 0.377 0.411 0.406 0.498 0.653 0.431 0.435 0.496 0.578
144 0.397 0.393 0.426 0.424 0.536 0.602 0.442 0.453 0.521 0.761
192 0.408 0.402 0.434 0.434 0.543 0.684 0.457 0.471 0.568 0.598
ETTh2 48 0.293 0.193 0.301 0.241 0.357 0.339 0.329 0.297 0.414 0.458
96 0.337 0.308 0.342 0.306 0.417 0.421 0.375 0.376 0.425 0.490
144 0.363 0.365 0.368 0.338 0.489 0.583 0.401 0.420 0.497 0.680
192 0.385 0.387 0.384 0.353 0.488 0.454 0.457 0.465 0.512 0.682
ETTm1 48 0.334 0.315 0.333 0.303 0.395 0.412 0.354 0.348 0.529 0.670
96 0.335 0.306 0.347 0.347 0.406 0.492 0.353 0.318 0.482 0.624
144 0.347 0.322 0.359 0.341 0.466 0.562 0.367 0.345 0.512 0.683
192 0.356 0.334 0.294 0.347 0.454 0.533 0.376 0.361 0.531 0.591
ETTm2 48 0.234 0.149 0.231 0.152 0.254 0.189 0.247 0.192 0.279 0.241
96 0.252 0.188 0.254 0.179 0.307 0.244 0.278 0.265 0.372 0.364
144 0.272 0.227 0.277 0.216 0.342 0.288 0.309 0.288 0.408 0.352
192 0.287 0.239 0.294 0.236 0.543 0.341 0.327 0.43 0.485 0.892
Exchange 48 0.139 0.045 0.149 0.046 0.187 0.073 0.159 0.067 0.205 0.124
96 0.196 0.083 0.211 0.092 0.294 0.159 0.247 0.168 0.778 0.409
144 0.244 0.113 0.265 0.141 0.375 0.292 0.272 0.31 0.68 0.671
192 0.288 0.180 0.329 0.212 0.464 0.494 0.354 0.403 0.979 0.544
Electricity 48 0.265 0.194 0.275 0.175 0.259 0.156 0.287 0.209 0.290 0.198
96 0.243 0.152 0.285 0.184 0.277 0.180 0.271 0.181 0.303 0.201
144 0.243 0.151 0.291 0.192 0.287 0.191 0.278 0.186 0.320 0.211
192 0.245 0.152 0.299 0.198 0.296 0.189 0.291 0.204 0.319 0.216
ILI 48 1.004 2.515 1.284 3.612 0.919 2.332 1.895 5.743 1.151 2.984
96 1.013 2.606 1.414 3.857 0.952 2.443 1.823 5.682 1.272 3.076
144 1.002 2.614 1.388 3.875 1.037 2.681 1.740 5.682 1.124 2.688
192 0.995 2.504 1.455 4.180 1.106 2.749 1.794 6.177 1.163 2.880
4.2. Comparative Experiment
We present the forecasting results in Table 1, where Red indicates the best result and Blue refers to the second best. Across Table 1, TimeSieve consistently outperformed the baseline models in most scenarios, particularly excelling in datasets with redundant features, such as ETTh1 and ETTh2. In the ETTh1 dataset, TimeSieve achieved a significant reduction in both MAE and MSE at all prediction horizons. At the 48-time step horizon in ETTh1, TimeSieve achieved an MAE of 0.361 and an MSE of 0.341, outperforming Koopa by approximately 6.2% in MAE and 6.3% in MSE. This consistent performance can be attributed to the effective integration of the Wavelet Decomposition Block (WDB) and the Information Filtering and Compression Block (IFCB), which together capture multi-scale features while filtering out redundant features. However, it did not achieve the best results in some experiments with a forecast window of 48. We believe this may be due to the significant short-term fluctuations in time series data, making it difficult for the model to distinguish meaningful signals from redundant features in shorter forecast windows. In contrast, longer windows (e.g., 96, 144, 192) provide more historical data, enabling the model to better capture long-term dependencies and seasonal patterns, and to differentiate redundant features from short-term features.
In datasets like ETTh2 and ETTm1, the performance improvements are also notable. TimeSieve shows a considerable advantage in short to medium-term predictions (48 to 144 steps). For instance, in ETTh2 at the 48-time step horizon, TimeSieve achieved an MAE of 0.293, improving upon Koopa’s MAE of 0.301 and LightT’s 0.329.
Interestingly, the Exchange dataset showed the most significant relative improvement. At the 48-time step horizon, TimeSieve achieved an MAE of 0.139 and an MSE of 0.045, outperforming Koopa by 6.7% in MAE and 2.2% in MSE. The superior performance on this dataset can be attributed to the effective redundant features filtering by the IFCB, which is crucial for financial data often characterized by high volatility.
However, in the ILI dataset, TimeSieve’s performance, while still competitive, showed less dramatic improvements compared to other datasets. For the 48-time step horizon, TimeSieve’s MAE was 1.004, compared to Autoformer’s 1.151. The relatively modest improvement can be explained by the inherent complexity and redundant features levels in the ILI data, which present significant challenges for all models. Despite this, TimeSieve’s ability to manage and slightly outperform in such a challenging dataset.
The varying performance across different datasets and prediction horizons (48, 96, 144, 192 steps) underscores the model’s versatility. TimeSieve consistently demonstrated the substantial performance gains at shorter horizons (48 and 96 steps), where the impact of redundant features and abrupt changes is more pronounced. This is particularly evident in the ETTh1 and ETTh2 datasets, where shorter-term predictions are crucial for timely decision-making. The integrated wavelet transform and information bottleneck approach effectively balance feature extraction and redundant feature compression, ensuring a certain level of reliable predictions even when data is limited.
In conclusion, TimeSieve’s ability to achieve state-of-the-art performance on 70% of the datasets tested highlights its superior predictive accuracy and generalization capability. A comparison of model residuals for TimeSieve, Koopa, LightTS, NSTformer, and Autoformer on the ETTh2 dataset with $H=48$ and $T=96$ is shown in Figure 3, visualizing the residuals between predicted and actual values.
4.3. Ablation Study
To validate the effectiveness of the modules used in our proposed model, we conducted ablation studies focusing on the wavelet transformation and the IB components. Further discussions on the selection of wavelet basis functions and the impact of $\mathcal{L}_{IB}$ loss weights can be found in Appendix Tables 6 and Tables 7.
Wavelet Transformation.
Table 2. Comparison of sequence decomposition methods with and without wavelet transformations (WDB & WRB and No Wavelet) on Exchange and ETTh2 datasets. We set the lookback length T = 2H.
Method No Wavelet WDB&WRB
[…]
Exchange 48 0.151 0.050 0.119 0.044
96 0.204 0.093 0.196 0.083
144 0.261 0.202 0.241 0.116
192 0.312 0.251 0.283 0.232
ETTh2 48 0.299 0.246 0.293 0.193
96 0.342 0.333 0.337 0.308
144 0.370 0.374 0.363 0.365
192 0.389 0.402 0.385 0.387
In Table 2, we compare the performance of our sequence decomposition methods with and without wavelet transformations(WDB&WRB and No Wavelet) on the Exchange and ETTh2 datasets. The results demonstrate that incorporating wavelet transformations significantly improves forecasting accuracy, as evidenced by lower MAE and MSE values across all forecasting horizons.
By leveraging wavelet transformations, our model can decompose the time series data into different frequency components, effectively compressing redundant features. This process enhances the model’s ability to capture the true underlying trends and patterns, leading to significant improvements in both MAE and MSE. For example, at the 96-step horizon on the Exchange dataset, the MAE decreases from 0.204 to 0.196 and the MSE decreases from 0.093 to 0.083. Similarly, on the ETTh2 dataset, the MAE decreases from 0.342 to 0.337 and the MSE decreases from 0.333 to 0.308.
The results presented in Table 2 validate the effectiveness of using wavelet transformations for time series decomposition. By leveraging the multi-scale nature of wavelets, our method can effectively capture both the coarse and fine-grained structures in the data, leading to more accurate and reliable forecasts.
Information Bottleneck
Table 3. Comparison of IFCB performance on Exchange and ETTh2 datasets under different conditions. $OnlyA$ applies IFCB only on the approximation coefficients after WDB, $OnlyD$ applies IFCB only on the detail coefficients after WRB, and ’TimeSieve’ applies IFCB on all coefficients.
Models TimeSieve $OnlyA$ $OnlyD$
[…]
Exchange 48 0.139 0.045 0.140 0.047 0.139 0.046
96 0.196 0.083 0.198 0.084 0.196 0.083
144 0.244 0.116 0.245 0.127 0.245 0.133
192 0.283 0.180 0.285 0.215 0.288 0.188
ETTh2 48 0.293 0.193 0.293 0.200 0.293 0.239
96 0.337 0.308 0.337 0.311 0.338 0.312
144 0.363 0.365 0.363 0.389 0.369 0.389
192 0.385 0.387 0.385 0.408 0.385 0.406
In Table 3, we present a comparison of our proposed TimeSieve method against two baseline methods, $OnlyA$ and $OnlyD$, on the Exchange and ETTh2 datasets. The results indicate that TimeSieve consistently outperforms the baseline methods across all forecasting horizons. However, the improvements are more pronounced in the Exchange dataset compared to the ETTh2 dataset.
The ETTh2 dataset contains fewer redundant features, making it relatively easier for traditional methods to capture the underlying patterns effectively. Consequently, the improvements from using TimeSieve are less pronounced. By leveraging the Information Filtering and Compression Block (IFCB), TimeSieve can effectively compress redundant features while retaining critical information. This process enhances the model’s ability to capture the true underlying trends and patterns, leading to significant improvements in both MAE and MSE. For example, at the 96-step horizon, TimeSieve records a MAE of 0.337 and an MSE of 0.308, outperforming both $OnlyA$ (MAE 0.337, MSE 0.311) and $OnlyD$ (MAE 0.338, MSE 0.312). In this context, $OnlyA$ refers to applying IFCB only on the approximation coefficients, and $OnlyD$ refers to applying IFCB only on the detail coefficients.
In contrast, the Exchange dataset, being a financial exchange rate dataset, exhibits relatively more redundant features compared to ETTh2. This makes the benefit from redundant features reduction provided by IFCB more pronounced. Consequently, the improvements in MAE for TimeSieve are relatively modest compared to the baseline methods. For instance, at the 192-step horizon, TimeSieve achieves a MAE of 0.283, only slightly better than $OnlyA$ (0.285) and $OnlyD$ (0.288). However, TimeSieve demonstrates more significant improvements in MSE. Since MSE emphasizes larger errors due to squaring the error terms, it better captures the overall reduction in error magnitude provided by IFCB. At the 192-step horizon, TimeSieve reduces the MSE to 0.180, compared to 0.215 for $OnlyA$ and 0.188 for $OnlyD$. This indicates that while individual large errors (affecting MAE) may still occur, the overall error distribution is improved with TimeSieve.
The results presented in Table 3 validate the effectiveness of using IFCB for time series decomposition. By leveraging IFCB, our method can effectively capture both the coarse and fine-grained structures in the data, leading to more accurate and reliable forecasts
The results presented in Table 3 highlight the superior performance of the TimeSieve method, demonstrating the effectiveness of the Information Filtering and Compression Block (IFCB) in reducing redundant features and enhancing model accuracy.
5. Conclusions
In this paper, we presented the TimeSieve framework for time series forecasting, emphasizing the decomposition and reconstruction of time series data through wavelet and information bottleneck techniques. Our approach integrates Wavelet Decomposition Block (WDB), Wavelet Reconstruction Block (WRB), and Information Fusion and Compression Block (IFCB) to enhance prediction accuracy.
We conducted extensive experiments on multiple benchmark datasets, demonstrating the superiority of TimeSieve in comparison to state-of-the-art methods. The results highlight the model’s capacity to handle varying forecast lengths and different types of time series data.
The proposed TimeSieve framework paves the way for future research in time series forecasting, offering a robust and scalable solution that can be applied to diverse domains such as finance, healthcare, and climate science. Future work will focus on further optimizing the model, exploring adaptive wavelet transforms, and extending the framework to multi-variate and multi-modal time series data.
Appendix A Markov Chain Theory
A.1. Basic Concepts
Markov chains are a powerful tool for describing the transitions between states in a system, widely used in various fields such as physics, biology, economics, and computer science. In deep learning and information theory, Markov chains can model the transmission of information and state dependencies. This document will introduce the basic concepts, properties, and applications of Markov chains in information theory.
A Markov chain is a sequence of random variables $\{X_{t}\}_{t\in T}$, where each random variable $X_{t}$ represents the state of the system at time $t$. The core property of a Markov chain is the Markov property, which states that the current state $X_{t}$ depends only on the previous state $X_{t-1}$ and not on any earlier states. Formally, this is expressed as:
$P(X_{t+1}=x_{t+1}\mid X_{t}=x_{t},X_{t-1}=x_{t-1},\ldots,X_{0}=x_{0})$
$\quad=P(X_{t+1}=x_{t+1}\mid X_{t}=x_{t})$
The state space $S$ is the set of all possible states of the Markov chain. For a finite state space, a Markov chain can be represented by a transition matrix $P$, where $P_{ij}$ denotes the probability of transitioning from state $i$ to state $j$:
$P=\begin{bmatrix}P_{11}&P_{12}&\cdots&P_{1n}\\ P_{21}&P_{22}&\cdots&P_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ P_{n1}&P_{n2}&\cdots&P_{nn}\end{bmatrix}$
The transition probabilities must satisfy the following conditions:
1. Non-negativity: $P_{ij}\geq 0$ 2. Normalization: $\sum_{j\in S}P_{ij}=1$
A probability vector $\pi$ is a stationary distribution of a Markov chain if it satisfies $\pi P=\pi$. For an irreducible and aperiodic Markov chain, the stationary distribution exists uniquely, and the state distribution converges to the stationary distribution regardless of the initial distribution:
$\lim_{t\to\infty}P(X_{t}=j)=\pi_{j}$
A.2. Applications in Information Theory
In information theory, Markov chains are commonly used to model the process of information transmission. In particular, in deep learning, hidden layers can be viewed as intermediate states in the information transmission process, where the relationship between input, hidden layer, and output can be modeled as a Markov chain. Assuming the input, hidden layer, and output are denoted as $X$, $Z$, and $Y$ respectively, we have the Markov property $X\rightarrow Z\rightarrow Y$, meaning:
$P(Y\mid Z,X)=P(Y\mid Z)$
Mutual information measures the mutual dependence between two random variables . For the above relationship, mutual information satisfies the following inequality:
$I(X;Z)\geq I(X;Y)$
This indicates that in the process of information transmission, the intermediate layer $Z$ retains more information about the input $X$ than the output $Y$, thereby better predicting the output $Y$.
Appendix B Mutual Information
B.1. Basic Concepts
Mutual Information (MI) is a fundamental concept in information theory that measures the amount of information obtained about one random variable through another random variable. It quantifies the reduction in uncertainty of one variable given knowledge of the other. In the context of deep learning, mutual information can help analyze the dependency and information flow between different layers of a neural network. This section will introduce the basic concepts, properties, and applications of mutual information.
Mutual information between two random variables $X$ and $Y$ is defined as the reduction in the uncertainty of $X$ due to the knowledge of $Y$. Formally, it is given by:
$I(X;Y)=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\log\left(\frac{p(x,y)}{p(x)p(y)}\right)$
where $p(x,y)$ is the joint probability distribution of $X$ and $Y$, and $p(x)$ and $p(y)$ are the marginal probability distributions of $X$ and $Y$, respectively.
Alternatively, mutual information can be expressed in terms of entropy:
$I(X;Y)=H(X)-H(X\mid Y)=H(Y)-H(Y\mid X)=H(X)+H(Y)-H(X,Y)$
where $H(X)$ and $H(Y)$ are the entropies of $X$ and $Y$, $H(X\mid Y)$ and $H(Y\mid X)$ are the conditional entropies, and $H(X,Y)$ is the joint entropy.
B.2. Properties of Mutual Information
Mutual information has several important properties that make it a useful measure in information theory:
  1. (1)
    Non-negativity: $I(X;Y)\geq 0$. Mutual information is always non-negative.
  2. (2)
    Symmetry: $I(X;Y)=I(Y;X)$. Mutual information is symmetric with respect to $X$ and $Y$.
  3. (3)
    Chain Rule: Mutual information satisfies a chain rule similar to entropy. For three random variables $X$, $Y$, and $Z$:
    $I(X;Y,Z)=I(X;Z)+I(X;Y\mid Z)$
  4. (4)
    Data Processing Inequality: If $X\rightarrow Y\rightarrow Z$ forms a Markov chain, then mutual information satisfies the inequality:
    $I(X;Z)\leq I(X;Y)$
B.3. Applications in Deep Learning
In deep learning, mutual information is used to understand and optimize the information flow through the layers of a neural network. Specifically, the Information Bottleneck (IB) method leverages mutual information to balance compression and prediction. The IB method aims to find a representation $Z$ that retains the most relevant information about the output $Y$ while minimizing the information about the input $X$. Formally, this is expressed as:
$\min\left[I(X;Z)-\beta I(Z;Y)\right]$
where $\beta$ is a trade-off parameter. This approach helps in understanding the trade-offs between compressing the input information and retaining predictive power.
B.4. Mutual Information Estimation
Estimating mutual information from data can be challenging due to the need to estimate the joint and marginal distributions. Various methods have been proposed for mutual information estimation, including:
  1. (1)
    Histogram-based methods: These methods discretize the data into bins and estimate probabilities from the frequency of occurrences.
  2. (2)
    Kernel density estimation: This non-parametric approach uses kernels to estimate the probability density functions.
  3. (3)
    Neural estimation methods: Recent advances use neural networks to approximate mutual information, such as MINE (Mutual Information Neural Estimation).
Appendix C Implementation Details
C.1. Baselines
C.1.1. Koopa (Liu et al., 2023a)
Koopa is a novel forecasting model designed to handle the intrinsic non-stationarity in real-world time series. Developed by researchers at Tsinghua University, Koopa leverages modern Koopman theory to address the complexities associated with changing temporal distributions. The model decomposes intricate non-stationary series into time-variant and time-invariant components using a Fourier Filter. Subsequently, it employs specialized Koopman Predictors to advance the dynamics of these components independently.
The innovative architecture of Koopa involves stackable blocks that learn hierarchical dynamics, facilitating a modular approach to time series forecasting. Each block integrates Koopman embedding and utilizes Koopman operators as linear representations of implicit transitions. For time-variant dynamics, the model calculates context-aware operators within the temporal neighborhood, ensuring adaptability to localized changes. This method enables Koopa to effectively extend the forecast horizon by incorporating incoming ground truth data.
C.1.2. NSTformer (Zeng et al., 2022)
The Nonstationary Transformer (NTSformer) is an advanced framework designed to address the challenges posed by non-stationary time series data in forecasting tasks. Developed by researchers at Tsinghua University, this model builds upon the conventional Transformer architecture, which excels at capturing temporal dependencies through its attention mechanism but struggles with non-stationary data where statistical properties and joint distributions change over time.
To mitigate the limitations of standard Transformers, the Nonstationary Transformer introduces two key modules: Series Stationarization and De-stationary Attention. Series Stationarization employs a normalization strategy to align the statistical properties of input time series, enhancing predictability by ensuring a more stable data distribution. This module normalizes the input series to a common scale, reducing the discrepancies caused by varying means and variances.
However, purely stationarizing the series can lead to the over-stationarization problem, where the model fails to capture crucial temporal dependencies intrinsic to non-stationary data. To address this, the De-stationary Attention module reintegrates the non-stationary characteristics into the attention mechanism. This module learns de-stationary factors that adjust the attention weights, preserving the unique temporal patterns of the original series.
C.1.3. LightTS (Campos et al., 2023)
LightTS is an innovative framework designed to perform lightweight time series classification, addressing the computational constraints often encountered in resource-limited environments such as edge devices. Developed by researchers from Aalborg University and several collaborating institutions, LightTS utilizes adaptive ensemble distillation to compress large, resource-intensive ensemble models into compact, efficient classifiers without compromising on accuracy.
The core of LightTS involves two primary innovations: adaptive ensemble distillation and Pareto optimal model selection. Adaptive ensemble distillation assigns dynamic weights to different base models, tailoring their contributions based on their classification capabilities, thereby distilling the ensemble knowledge into a single lightweight model. This approach ensures that the most relevant information is captured efficiently, leading to a high-performing student model. Furthermore, LightTS employs a multi-objective Bayesian optimization method to identify Pareto optimal settings that balance model accuracy and size, enabling users to select the most accurate lightweight models that fit within specific storage constraints.
C.1.4. Autoformer (Wu et al., 2021)
Autoformer is a framework designed to tackle the complexities of long-term time series forecasting. Developed by researchers at Tsinghua University, Autoformer innovatively combines decomposition architecture with an Auto-Correlation mechanism to enhance the model’s ability to handle intricate temporal patterns. Traditional Transformer-based models often struggle with long-term dependencies due to the prohibitive computational costs and the obscured nature of temporal dependencies in extended forecasting horizons. Autoformer addresses these challenges by embedding series decomposition blocks within the deep learning architecture, allowing for progressive extraction and refinement of long-term trends and seasonal components.
The Auto-Correlation mechanism is a key innovation of Autoformer, inspired by stochastic process theory. This mechanism leverages the inherent periodicity in time series data to discover dependencies and aggregate similar sub-series efficiently. By replacing the self-attention mechanism with Auto-Correlation, Autoformer achieves a significant reduction in computational complexity, scaling as O(L logL) for a series of length L. This enables more efficient and accurate long-term forecasting compared to previous models.
C.2. Datasets
Exchange Rate Dataset: The Exchange Rate dataset consists of the daily exchange rates of eight different countries, spanning from 1990 to 2016. This dataset is commonly used to study and forecast currency fluctuations, providing valuable insights for economic planning and financial analysis. The data includes exchange rates for currencies such as USD, GBP, EUR, JPY, and others.
ETT (Electricity Transformer Temperature) Dataset: The ETT dataset includes data collected from electricity transformers, capturing both load and oil temperature measurements at 15-minute intervals over a period from July 2016 to July 2018. This dataset is particularly useful for energy consumption forecasting and transformer health monitoring. The dataset is split into ETTh1, ETTh2, ETTm1, and ETTm2 subsets, representing different transformer units and sampling frequencies (hourly and 15-minutely).
Electricity Consumption Dataset: The Electricity Consumption dataset contains hourly electricity usage data of 321 clients from 2012 to 2014. This dataset is utilized to predict future electricity demand and helps in managing energy resources efficiently. The data captures variations in electricity usage patterns, which are influenced by factors such as weather, economic conditions, and consumer behavior.
National Illness Dataset: The National Illness dataset, also known as the Influenza-like Illness (ILI) dataset, comprises weekly recorded data from the Centers for Disease Control and Prevention (CDC) of the United States. It covers the period from 2002 to 2021 and includes the ratio of patients exhibiting influenza-like symptoms to the total number of patients seen. This dataset is essential for modeling and forecasting disease outbreaks, aiding in public health planning and response strategies.
Table 4. Summary of Datasets Used in Time Series Forecasting
Dataset Time Points Frequency Variables
ETTh1 17420 Hour 7
ETTh2 17420 Hour 7
ETTm1 69680 15 Min 7
ETTm2 69680 15 Min 7
Exchange Rate 7588 Day 8
Electricity 26304 Hour 321
National Illness 966 Week 7
C.3. Definition of Metric
Mean Absolute Error (MAE): The Mean Absolute Error is a metric used to measure the average magnitude of errors between predicted values and actual values. It is defined as follows:
(18) $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|$
where $y_{i}$ represents the actual value, $\hat{y}_{i}$ represents the predicted value, and $n$ is the number of observations.
Mean Squared Error (MSE): The Mean Squared Error is another metric used to measure the average of the squares of the errors between predicted values and actual values. It is defined as follows:
(19) $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2}$
where $y_{i}$ represents the actual value, $\hat{y}_{i}$ represents the predicted value, and $n$ is the number of observations.
C.4. Experimental Setup
Component Description
GPU NVIDIA GeForce RTX 3080
CPU Intel(R) Xeon(R) E5-2686 v4
Epochs 10
Evaluation Metrics MAE and MSE
Batch Size 32
Optimizer Adam
Learning Rate 0.0001
Framework PyTorch 2.0.1
CUDA Version CUDA 11.8
Table 5. Experimental Setup: Summary of Hardware and Software Configurations
Table 6. Comparison of different $\mathcal{L}_{IB}$ loss weights on Exchange rate dataset. The lookback length is set to $T=2H$.
IB Weight 0.000001 0.00001 0.001 0.01 0.1
[…]
48 0.139374 0.050472 0.139033 0.045016 0.139010 0.044790 0.138870 0.044050 0.139096 0.043964
96 0.194557 0.110181 0.196048 0.083312 0.195420 0.084286 0.195818 0.085749 0.194953 0.085362
144 0.249113 0.139211 0.243771 0.133058 0.241323 0.116296 0.241018 0.153478 0.240288 0.152741
192 0.300818 0.1886299 0.287500 0.1804375 0.283648 0.1815776 0.283406 0.1777823 0.289262 0.1804006
Table 7. Comparison of different wavelet basis functions on Exchange rate dataset.The lookback length is set to $T=2H$.
Metric HAAR SYM2 DB1
[…]
48 0.1392049 0.0477886 0.1419515 0.0450065 0.1390332 0.0450161
96 0.1944531 0.1024204 0.1944787 0.0895267 0.1960484 0.0833123
144 0.2446094 0.1230493 0.2429469 0.1325571 0.2437714 0.1330583
192 0.3007113 0.2840106 0.2913505 0.1804375 0.2875000 0.1875995
Figure 3. Comparison of model residuals for TimeSieve, Koopa, LightTS, NSTformer, and Autoformer on the ETTh2 dataset with $H=48$ and $T=96$. The vertical axis represents the absolute difference between predicted and actual values.